# Bandit and Delayed Feedback in Online Structured Prediction

### v1

@@ 2.2 Fenchel–Young Loss
We use the Fenchel–Young loss (Blondel et al., 2020) as the surrogate loss, which subsumes many representative surrogate losses, such as logistic loss, Conditional Random Field (CRF) loss (Lafferty et al., 2001), and SparseMAP (Niculae et al., 2018). See Blondel et al. (2020, Table 1) for more examples.
Definition 2.2 (Blondel et al. 2020, Fenchel–Young loss).
Let $\Omega:\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a regularization function with $\mathcal{Y}\subset\operatorname{dom}(\Omega)$. The Fenchel–Young loss generated by $\Omega$, denoted by $S_{\Omega}:\operatorname{dom}(\Omega^{\ast})\times\operatorname{dom}(\Omega)\rightarrow\mathbb{R}_{\geq 0}$, is defined as
$S_{\Omega}(\bm{\theta};\bm{y})\coloneqq\Omega^{\ast}(\bm{\theta})+\Omega(\bm{y})-\langle\bm{\theta},\bm{y}\rangle.$
The Fenchel–Young loss has the following properties, which will be useful in the subsequent discussion:
Proposition 2.3 (Blondel et al. 2020, Propositions 2 and 3 and Sakaue et al. 2024, Proposition 3).
@@
$\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ $\coloneqq\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Omega(\bm{y})\>:\>\bm{y}\in\mathbb{R}^{d}\}$
$=\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Psi(\bm{y})\>:\>\bm{y}\in\operatorname{conv}(\mathcal{Y})\}.$
Then, for any $\bm{y}\in\mathcal{Y}$, $S_{\Omega}(\bm{\theta},\bm{y})$ is differentiable with respect to $\bm{\theta}$, and it satisfies $\nabla S_{\Omega}(\bm{\theta};\bm{y})=\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{y}.$ Furthermore, it holds that $S_{\Omega}(\bm{\theta};\bm{y})\geq\frac{\lambda}{2}\|\bm{y}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|^{2}.$
@@
Indeed, from Proposition 2.3 and Assumption 2.1, we have $\|\nabla S_{t}(\bm{W}_{t})\|_{\mathrm{F}}^{2}=\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|_{2}^{2}\|\bm{x}_{t}\|^{2}\leq C_{x}^{2}\kappa^{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|^{2}\leq\frac{2C_{x}^{2}\kappa^{2}}{\lambda}S_{t}(\bm{W}_{t})$, where we used $\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$ and $\|\cdot\|_{2}\leq\kappa\|\cdot\|$. Thus, (2) holds with $b=\frac{2C_{x}^{2}\kappa^{2}}{\lambda}$. Below, let $L_{t}(\bm{y})\coloneqq L(\bm{y};\bm{y}_{t})$, $\bm{G}_{t}\coloneqq\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$, and $L\coloneqq\max_{\bm{W}\in\mathcal{W}}\|\nabla S_{t}(\bm{W})\|_{\mathrm{F}}$.
@@ 2.3 Examples of Structured Prediction
We present several special cases of structured prediction along with specific parameter values introduced so far; see Sakaue et al. (2024, Section 2.3) for further details.
@@
Let $\mathcal{Y}=\{\mathbf{e}_{1},\dots,\mathbf{e}_{d}\}$ and $\|\cdot\|=\|\cdot\|_{1}$. When using the 0-1 loss, $L(\bm{y}^{\prime};\bm{y})=\mathds{1}[\bm{y}^{\prime}\neq\bm{y}]$, the parameters in Assumption 2.1 are $\nu=2$ and $\gamma=\frac{1}{2}$. The logistic surrogate loss is a Fenchel–Young loss $S_{\Omega}$ generated by the entropy regularization function $\Omega=\mathsf{H}^{s}+I_{\Delta_{d}}$ (up to a constant factor), where $\mathsf{H}^{s}(\bm{y})\coloneqq-\sum_{i=1}^{d}y_{i}\log y_{i}$ and $\Delta_{d}$ is the $(d-1)$-dimensional probability simplex. Since $\Omega$ is a $1$-strongly convex function with respect to $\|\cdot\|_{1}$, we have $\lambda=1$.
@@
Let $\mathcal{Y}=\{0,1\}^{d}$ and $\|\cdot\|=\|\cdot\|_{2}$. When using the Hamming loss as the target loss function $L(\bm{y}^{\prime};\bm{y})=\frac{1}{d}\sum_{i=1}^{d}\mathds{1}[y_{i}^{\prime}\neq y_{i}]$, Assumption 2.1 is satisfied with $\nu=1$ and $\gamma=\frac{1}{d}$. The SparseMAP surrogate loss $S_{\Omega}(\bm{\theta},\bm{y})=\frac{1}{2}\|\bm{y}-\bm{\theta}\|_{2}^{2}-\frac{1}{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{\theta}\|_{2}^{2}$ is a Fenchel–Young loss generated by $\Omega=\frac{1}{2}\|\cdot\|^{2}+I_{\operatorname{conv}(\mathcal{Y})}$. Since $\Omega$ is 1-strongly convex with respect to $\|\cdot\|_{2}$, we have $\lambda=1$.
@@
We consider predicting the ranking of $m$ items. Let $\|\cdot\|=\|\cdot\|_{1}$, $d=m^{2}$, and $\mathcal{Y}\subset\{0,1\}^{d}$ be the set of all vectors representing $m\times m$ permutation matrices. We use the target loss function that counts mismatches, $L(\bm{y}^{\prime};\bm{y})=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[y_{i,j_{i}}^{\prime}\neq y_{i,j_{i}}]$, where $j_{i}\in[m]$ is a unique index with $y_{ij_{i}}=1$ for each $i\in[m]$. In this case, the parameters in Assumption 2.1 satisfy $\nu=4$ and $\gamma=\frac{1}{2m}$. We use a surrogate loss given by $S_{\Omega}(\bm{\theta};\bm{y})=\langle\bm{\theta},\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{y}\rangle+\frac{1}{\zeta}\mathsf{H}^{s}(\widehat{\bm{y}}_{\Omega}(\bm{\theta}))$, where $\Omega=-\frac{1}{\zeta}\mathsf{H}^{s}+I_{\operatorname{conv}(\mathcal{Y})}$ and $\zeta$ is a parameter controlling the regularization strength. The first term in $S_{\Omega}$ measures the affinity between $\bm{\theta}$ and $\bm{y}$, while the second term evaluates the uncertainty of $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$. Since $\Omega$ is $\frac{1}{m\zeta}$-strongly convex, we have $\lambda=\frac{1}{m\zeta}$.
@@ 2.4 Structured Encoding Loss Function (SELF)
Here we introduce a common class of target loss functions, called the Structured Encoding Loss Function (SELF). A target loss function is SELF if it can be expressed as
@@
where $\bm{b}\in\mathbb{R}^{d}$ is a constant vector, $\bm{V}\in\mathbb{R}^{d\times d}$ is a constant matrix, and $c:\mathcal{Y}\to\mathbb{R}$ is a function. The following loss examples, taken from Blondel (2019, Appendix A), belong to the SELF class:
@@
  • Multilabel classification: the Hamming loss, $L(\bm{y}^{\prime};\bm{y})=\frac{1}{d}\sum_{i=1}^{d}\mathds{1}[y^{\prime}_{t,i}\neq y_{i}]$, is a SELF with $\bm{V}=-\frac{2}{d}\bm{I}$, $\bm{b}=\frac{\bm{1}}{d}$, and $c(\bm{y})=\frac{1}{d}\langle\bm{y},\bm{1}\rangle$, where the last factor is constant if the number of correct labels is fixed.
@@
Following Sakaue et al. (2024), this study assumes that the target loss function $L$ is a SELF.
@@ 2.5 Randomized Decoding
2:$\widehat{\bm{y}}_{\Omega}(\bm{\theta})\leftarrow\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Psi(\bm{y})\>:\>\bm{y}\in\operatorname{conv}(\mathcal{Y})\}$
3:$\bm{y}^{\ast}\leftarrow\operatorname*{arg\,min}\{\|\bm{y}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|\>:\>\bm{y}\in\mathcal{Y}\}$
4:$\Delta^{\ast}\leftarrow\|\bm{y}^{\ast}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|,\>p\leftarrow\min\{1,2\Delta^{\ast}/\nu\}$
5:Sample $Z\sim\mathrm{Ber}(p)$
6:if $Z=0$ then $\bm{\widehat{y}}\leftarrow\bm{y}^{\ast}$
7:if $Z=1$ then $\bm{\widehat{y}}\leftarrow\bm{\widetilde{y}}$ where $\bm{\widetilde{y}}$ is randomly drawn from $\mathcal{Y}$ so that $\mathbb{E}\left[\bm{\widetilde{y}}|Z=1\right]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$
@@
The procedure of converting the estimated score $\bm{\theta}$ into a structured output $\bm{\widehat{y}}$ is called decoding. For this, we employ randomized decoding (Sakaue et al., 2024), which plays an essential role particularly in deriving an upper bound independent of the output set size $K=\lvert\mathcal{Y}\rvert$ in Section 3.4. The randomized decoding (Algorithm 1) selects either the closest $\bm{y}^{*}\in\mathcal{Y}$ to $\widehat{\bm{y}}_{\Omega}(\bm{\theta})\in\operatorname{conv}(\mathcal{Y})$ or a random $\widetilde{\bm{y}}\in\mathcal{Y}$ satisfying $\mathbb{E}[\widetilde{\bm{y}}\mid Z=1]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$, where $Z$ follows a Bernoulli distribution with a parameter $p$. Intuitively, the parameter $p$ is chosen so that if $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ is close to $\bm{y}^{*}$, the decoding is more likely to return $\bm{y}^{*}$; otherwise, it is more likely to return $\widetilde{\bm{y}}$, reflecting uncertainty. An important property satisfied by the randomized decoding is the following lemma, which we use in the subsequent analysis:
Lemma 2.4 (Sakaue et al. 2024, Lemma 4).
For any $(\bm{\theta},\bm{y})\in\mathbb{R}^{d}\times\mathcal{Y}$, the randomized decoding $\phi_{\Omega}$ satisfies
$\mathbb{E}[L(\phi_{\Omega}(\bm{\theta});\bm{y})]\leq\frac{4\gamma}{\lambda\nu}S_{\Omega}(\bm{\theta};\bm{y}),$
where the expectation is taken with respect to the internal randomness of the randomized decoding.
@@ 3 Bandit Feedback
In this section, we present two online structured prediction algorithms in the bandit feedback setup and analyze their regret. Our results here are mostly special cases of the regret bounds obtained when handling bandit and delayed feedback (Section 5). Nevertheless, by focusing on the case without delay, we provide a clearer exposition of the core ideas.
@@ 3.1 Randomized Decoding with Uniform Exploration
2:Sample $X\sim\mathrm{Ber}(q)$
3:if $X=0$ then $\bm{\widehat{y}}\leftarrow\phi_{\Omega}(\bm{\theta})$
4:if $X=1$ then Sample $\bm{y}^{\ast}\sim\mathrm{Unif}(\mathcal{Y})$ and $\bm{\widehat{y}}\leftarrow\bm{y}^{\ast}$
@@
Here, we discuss the properties of the decoding function, Randomized Decoding with Uniform Exploration (RDUE), which will be used in both algorithms to convert scores into outputs. As discussed in Section 2.5, in online structured prediction with full information feedback, the randomized decoding (Algorithm 1) was introduced as a decoding function (Sakaue et al., 2024). However, naively applying the randomized decoding does not lead to a satisfactory regret bound under bandit feedback. We extend the framework of the randomized decoding to handle bandit feedback effectively.
RDUE (Algorithm 2) is a procedure that, with probability $q\in[0,1]$, selects $\widehat{\bm{y}}$ uniformly at random from $\mathcal{Y}$, and with probability $1-q$, selects the output of the randomized decoding. Using RDUE, we define $p_{t}(\bm{y})$ as the probability that $\widehat{\bm{y}}_{t}$ coincides with $\bm{y}$ at round $t$. Note that for any $\bm{y}\in\mathcal{Y}$, it holds that $p_{t}(\bm{y})\geq\frac{q}{K}.$ Furthermore, similar to the property of the randomized decoding in Lemma 2.4, RDUE satisfies the following property:
@@
For any $(\bm{\theta},\bm{y})\in\mathbb{R}^{d}\times\mathcal{Y}$, RDUE $\psi_{\Omega}$ satisfies
$\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]\leq\frac{4\gamma}{\lambda\nu}(1-q)S_{\Omega}(\bm{\theta};\bm{y})+q\frac{K-1}{K},$
where the expectation is taken with respect to the internal randomness of RDUE.
@@
Since the output of the randomized decoding is chosen with probability $1-q$, and a uniformly random output is chosen with probability $q$, we have $\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]=(1-q)\mathbb{E}\left[L(\phi_{\Omega}(\bm{\theta});\bm{y})\right]+q\frac{K-1}{K},$ where we used $L(\cdot;\cdot)\leq 1$ and $\phi_{\Omega}$ is the randomized decoding. Hence, combining this with Lemma 2.4, we obtain the desired bound. ∎
Additionally, this section makes the following assumption:
@@
There exists $a\in(0,1)$ such that
$\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t})+q.$
Here, $\mathbb{E}_{t}\left[{\cdot}\right]$ denotes the conditional expectation given the random variables $\widehat{\bm{y}}_{1},\dots,\widehat{\bm{y}}_{t-1}$.
This assumption can be satisfied by using RDUE and letting $a\leq 1-\frac{4\gamma}{\lambda\nu}(1-q)$ if $\lambda>\frac{4\gamma}{\nu}(1-q)$, according to Lemma 3.1. In what follows, let $a=1-\frac{4\gamma}{\lambda\nu}$. Note that $\lambda\geq\frac{4\gamma}{\nu}$ holds in the cases of multiclass classification, multilabel classification, and ranking, as discussed in Section 2.3.

### v2

@@ 2.2 Fenchel–Young loss
We use the Fenchel–Young loss [6] as the surrogate loss, which subsumes many representative surrogate losses, such as the logistic loss, Conditional Random Field (CRF) loss [25], and SparseMAP [30]. See [6, Table 1] for more examples.
Definition 2.2 ([6, Fenchel–Young loss]).
Let $\Omega\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a regularization function with $\mathcal{Y}\subset\operatorname{dom}(\Omega)$. The Fenchel–Young loss generated by $\Omega$, denoted by $S_{\Omega}\colon\operatorname{dom}(\Omega^{\ast})\times\operatorname{dom}(\Omega)\rightarrow\mathbb{R}_{\geq 0}$, is defined as $S_{\Omega}(\bm{\theta};\bm{y})\coloneqq\Omega^{\ast}(\bm{\theta})+\Omega(\bm{y})-\langle\bm{\theta},\bm{y}\rangle.$
The Fenchel–Young loss has the following useful properties:
Proposition 2.3 ([6, Propositions 2 and 3] and [35, Proposition 3]).
@@
Indeed, from Proposition 2.3, we have $\|\nabla S_{t}(\bm{W}_{t})\|_{\mathrm{F}}^{2}=\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|_{2}^{2}\|\bm{x}_{t}\|_{2}^{2}\leq C_{x}^{2}\kappa^{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|^{2}\leq\frac{2C_{x}^{2}\kappa^{2}}{\lambda}S_{t}(\bm{W}_{t})$, where we used $\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$ and $\|\cdot\|_{2}\leq\kappa\|\cdot\|$. Thus, (1) holds with $b=\frac{2C_{x}^{2}\kappa^{2}}{\lambda}$. Below, let $L_{t}(\bm{y})\coloneqq L(\bm{y};\bm{y}_{t})$ and $\bm{G}_{t}\coloneqq\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$.
@@ 2.3 Examples of structured prediction
We present several instances of structured prediction along with specific parameter values introduced so far; see [35, Section 2.3] for further details.
@@
Let $\mathcal{Y}=\{\mathbf{e}_{1},\dots,\mathbf{e}_{d}\}$ and $\|\cdot\|=\|\cdot\|_{1}$. As the 0-1 loss satisfies $L(\bm{y};\mathbf{e}_{i})=\mathds{1}[\bm{y}\neq\bm{y}]=\sum_{j\neq i}y_{j}=\frac{1}{2}(1-y_{i}+\sum_{j\neq i}y_{j})=\frac{1}{2}\|\mathbf{e}_{i}-\bm{y}\|_{1}$, we have $\gamma=\frac{1}{2}$ in Assumption 2.1. Also, $\nu=2$ holds since $\|\mathbf{e}_{i}-\mathbf{e}_{j}\|_{1}=2$ if $i\neq j$. The logistic surrogate loss is a Fenchel–Young loss $S_{\Omega}$ generated by the negative Shannon entropy $\Omega=\mathsf{H}^{s}+I_{\Delta_{d}}$ (up to a constant factor originating from the base of $\log$), where $\mathsf{H}^{s}(\bm{y})\coloneqq-\sum_{i=1}^{d}y_{i}\log y_{i}$ and $\Delta_{d}$ is the $(d-1)$-dimensional probability simplex. Since $\Omega$ is a $1$-strongly convex function w.r.t. $\|\cdot\|_{1}$ on $\Delta_{d}$, we have $\lambda=1$.
@@
Let $\mathcal{Y}=\{0,1\}^{d}$ and $\|\cdot\|=\|\cdot\|_{2}$. When using the Hamming loss as the target loss function $L(\bm{y}^{\prime};\bm{y})=\frac{1}{d}\sum_{i=1}^{d}\mathds{1}[y_{i}^{\prime}\neq y_{i}]$, Assumption 2.1 is satisfied with $\nu=1$ and $\gamma=\frac{1}{d}$. The SparseMAP surrogate loss $S_{\Omega}(\bm{\theta},\bm{y})=\frac{1}{2}\|\bm{y}-\bm{\theta}\|_{2}^{2}-\frac{1}{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{\theta}\|_{2}^{2}$ is a Fenchel–Young loss generated by $\Omega=\frac{1}{2}\|\cdot\|^{2}+I_{\operatorname{conv}(\mathcal{Y})}$. Since $\Omega$ is 1-strongly convex w.r.t. $\|\cdot\|_{2}$, we have $\lambda=1$.
@@
We consider predicting the ranking of $m$ items. Let $\|\cdot\|=\|\cdot\|_{1}$, $d=m^{2}$, and $\mathcal{Y}\subset\{0,1\}^{d}$ be the set of all vectors representing $m\!\times\!m$ permutation matrices. We use the target loss function that counts mismatches, $L(\bm{y}^{\prime};\bm{y})=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[y_{i,j_{i}}^{\prime}\!\neq\!y_{i,j_{i}}]$, where $j_{i}\in[m]$ is a unique index with $y_{ij_{i}}=1$ for each $i\in[m]$. In this case, Assumption 2.1 is satisfied with $\nu=4$ and $\gamma=\frac{1}{2m}$. We use a surrogate loss given by $S_{\Omega}(\bm{\theta};\bm{y})\!=\!\langle\bm{\theta},\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{y}\rangle\!+\!\frac{1}{\zeta}\mathsf{H}^{s}(\widehat{\bm{y}}_{\Omega}(\bm{\theta}))$, where $\Omega\!=\!-\frac{1}{\zeta}\mathsf{H}^{s}\!+\!I_{\operatorname{conv}(\mathcal{Y})}$ and $\zeta$ controls the regularization strength. The first term in $S_{\Omega}$ measures the affinity between $\bm{\theta}$ and $\bm{y}$, while the second term evaluates the uncertainty of $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$. Since $\Omega$ is $\frac{1}{m\zeta}$-strongly convex, we have $\lambda=\frac{1}{m\zeta}$.
@@ 2.4 Structured encoding loss function (SELF)
We introduce a general class of target loss functions, called the (generalized) Structured Encoding Loss Function (SELF) [10, 11, 5]. A target loss function is SELF if it can be expressed as
@@
where $\bm{b}\in\mathbb{R}^{d}$ is a constant vector, $\bm{V}\in\mathbb{R}^{d\times d}$ is a constant matrix, and $c\colon\mathcal{Y}\to\mathbb{R}$ is a function. The following examples of target losses, taken from [5, Appendix A], belong to the SELF class:
@@
  • Multilabel classification: the Hamming loss, $L(\bm{y}^{\prime};\bm{y})\!=\!\frac{1}{d}\sum_{i=1}^{d}\mathds{1}[y^{\prime}_{t,i}\!\neq\!y_{i}]$, is a SELF with $\bm{V}\!=\!-\frac{2}{d}\bm{I}$, $\bm{b}\!=\!\frac{1}{d}\bm{1}$, and $c(\bm{y})\!=\!\frac{1}{d}\langle\bm{y},\bm{1}\rangle$, where $c(\bm{y})$ is constant if the number of correct labels is fixed.
@@
Following the work by Sakaue et al. [35], we assume that the target loss function $L$ is a SELF.
@@ 2.5 Randomized decoding
2:Compute $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ defined in Proposition 2.3.
3:$\bm{y}^{\ast}\leftarrow\operatorname*{arg\,min}\{\|\bm{y}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|\>:\>\bm{y}\in\mathcal{Y}\}$.
4:$\Delta^{\ast}\!\leftarrow\!\|\bm{y}^{\ast}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|,\>p\!\leftarrow\!\min\{1,2\Delta^{\ast}/\nu\}$.
5:Sample $Z\sim\mathrm{Ber}(p)$.
6:if $Z=0$ then $\bm{\widehat{y}}\leftarrow\bm{y}^{\ast}$.
7:if $Z=1$ then $\bm{\widehat{y}}\leftarrow\bm{\widetilde{y}}$ where $\bm{\widetilde{y}}$ is randomly drawn from $\mathcal{Y}$ so that $\mathbb{E}\left[\bm{\widetilde{y}}|Z=1\right]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$.
@@
We employ the randomized decoding [35], which plays an essential role, particularly in deriving an upper bound independent of the output set size $K=\lvert\mathcal{Y}\rvert$ in Section 3.4. The randomized decoding (Algorithm 1) returns either the closest $\bm{y}^{*}\in\mathcal{Y}$ to $\widehat{\bm{y}}_{\Omega}(\bm{\theta})\in\operatorname{conv}(\mathcal{Y})$ (see Proposition 2.3) or a random $\widetilde{\bm{y}}\in\mathcal{Y}$ satisfying $\mathbb{E}[\widetilde{\bm{y}}\mid Z=1]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$, where $Z$ follows the Bernoulli distribution with a parameter $p$. Intuitively, the parameter $p$ is chosen so that if $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ is close to $\bm{y}^{*}$, the decoding more likely returns $\bm{y}^{*}$; otherwise, it more likely returns $\widetilde{\bm{y}}$, reflecting uncertainty. An crucial property of the randomized decoding is the following lemma, which we use in the subsequent analysis:
Lemma 2.4 ([35, Lemma 4]).
For any $(\bm{\theta},\bm{y})\!\!\in\!\mathbb{R}^{d}\times\mathcal{Y}$, the randomized decoding $\phi_{\Omega}$ satisfies $\mathbb{E}[L(\phi_{\Omega}(\bm{\theta});\bm{y})]\!\leq\frac{4\gamma}{\lambda\nu}S_{\Omega}(\bm{\theta};\bm{y})$, where the expectation is taken w.r.t. the randomness of $\phi_{\Omega}$.
@@ 3 Bandit feedback
In this section, we present two OSP algorithms for the bandit feedback setting and analyze their surrogate regret. Our results can be extended to handle bandit and delayed feedback; see Section 5. Here, we focus on the simpler case without delay to provide a clearer exposition of our core ideas.
@@ 3.1 Randomized decoding with uniform exploration
2:Sample $X\sim\mathrm{Ber}(q)$.
3:if $X=0$ then ${\widehat{\bm{y}}}\leftarrow\phi_{\Omega}(\bm{\theta})$.
4:if $X=1$ then Sample $\bm{y}^{\ast}$ from $\mathcal{Y}$ uniformly at random and $\widehat{\bm{y}}\!\leftarrow\!\bm{y}^{\ast}$.
@@
We discuss the properties of our decoding function, called Randomized Decoding with Uniform Exploration (RDUE), which will be used in subsequent sections. As discussed in Section 2.5, the randomized decoding (Algorithm 1) was introduced as a decoding function [35] for OSP with full-information feedback. However, naively applying it does not lead to a desired bound under bandit feedback due to the lack of exploration. We extend the randomized-decoding framework to handle bandit feedback effectively.
RDUE (Algorithm 2) is a procedure that, with probability $q\in[0,1]$, selects $\widehat{\bm{y}}$ uniformly at random from $\mathcal{Y}$, and with probability $1-q$, selects the output of the randomized decoding. Let $p_{t}(\bm{y})$ be the probability that the output of RDUE, $\widehat{\bm{y}}_{t}$, coincides with $\bm{y}$ at round $t$. Note that for any $\bm{y}\in\mathcal{Y}$, it holds that $p_{t}(\bm{y})\geq\frac{q}{K}$ thanks to the uniform exploration. Furthermore, similar to the property of the randomized decoding in Lemma 2.4, RDUE satisfies the following property:
@@
For any $(\bm{\theta},\bm{y})\in\mathbb{R}^{d}\times\mathcal{Y}$, RDUE $\psi_{\Omega}$ satisfies $\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]\leq\frac{4\gamma}{\lambda\nu}(1-q)S_{\Omega}(\bm{\theta};\bm{y})+q\frac{K-1}{K},$ where the expectation is taken w.r.t. the internal randomness of RDUE.
@@
With probability $1-q$, the randomized decoding is used; otherwise, a uniformly random output is chosen. Thus, $\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]=(1-q)\mathbb{E}\left[L(\phi_{\Omega}(\bm{\theta});\bm{y})\right]+q\frac{K-1}{K}$ holds, where $\phi_{\Omega}$ is the randomized decoding and we used $L(\cdot;\cdot)\leq 1$. Combining this with Lemma 2.4 completes the proof.∎
Based on this lemma, we make the following assumption for convenience:
@@
There exists $a\in(0,1)$ such that $\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t})+q.$ Here, $\mathbb{E}_{t}\left[{\cdot}\right]$ denotes the conditional expectation given the past outputs, $\widehat{\bm{y}}_{1},\dots,\widehat{\bm{y}}_{t-1}$.
This assumption can be satisfied by using RDUE for $a\leq 1-\frac{4\gamma}{\lambda\nu}(1-q)$ if $\lambda>\frac{4\gamma}{\nu}(1-q)$, due to Lemma 3.1. In what follows, let $a=1-\frac{4\gamma}{\lambda\nu}$. Note that $\lambda>\frac{4\gamma}{\nu}\geq\frac{4\gamma}{\nu}(1-q)$ holds in the cases of multiclass classification, multilabel classification, and ranking (see Section 2.3 and [35] for details).